OpenStudy (anonymous):

Let m and n be two integers and p a positive integer, show that \[\left(m-\sqrt{m^2-4 n}\right)^p+\left(\sqrt{m^2-4 n}+m\right)^p \] is an integer divisible by \[ 2^p\]

OpenStudy (kainui):

Every day it seems like you ask a similar question, but they get harder and harder haha. I am going to try to solve it this time.

ganeshie8 (ganeshie8):

this question is equivalent to the previous two questions ;p

ganeshie8 (ganeshie8):

\[x^2 - mx + n = 0\] \[x = \frac{m \pm \sqrt{m^2-4n}}{2}\] prove that \(\alpha^p + \beta^p\) is an integer

OpenStudy (kainui):

Essentially it looks like these are just two roots of a quadratic equation. Naturally solving them will bring out 1/2 and you can simply raise both of these to arbitrary powers of p and the two will come along for the ride.

ganeshie8 (ganeshie8):

yes we still need to prove it though if i remember, the recent proof we had was for a simpler quadratic x^2 + x = 1

ganeshie8 (ganeshie8):

I'll try induction on \(p\) : \(F(p) = \left(m-\sqrt{m^2-4 n}\right)^p+\left(\sqrt{m^2-4 n}+m\right)^p\) is divisible by \(2^p\) \(F(1) = 2^1m \) \(F(2) = 2^2(m^2-2n)\) the statement holds for \(F(1)\) and \(F(2)\) suppose \(F(k)\) and \(F(k-1)\) are divisible by \(2^k\) and \(2^{k-1}\) then we have \(F(k) = 2^k x\) and \(F(k-1) = 2^{k-1}y\) for some integers \(x\) and \(y\) \[\begin{align}F(k+1) &= \left(m-\sqrt{m^2-4 n}\right)^{k+1}+\left(\sqrt{m^2-4 n}+m\right)^{k+1}\\~\\ &= 2m\times F(k) - 4n\times F(k-1)~~~\color{red}{\star}\\~\\ &= 2m\times 2^k x - 4n\times 2^{k-1}y\\~\\ &= 2^{k+1}(mx - ny) \end{align}\] \(\blacksquare\) \(\color{red}{*} : ~ a^{k+1}+b^{k+1} = (a+b)(a^k+b^k) - ab(a^{k-1}+b^{k-1})\)

OpenStudy (anonymous):

\[r = \frac {1} {2} \left (m - \sqrt {m^2 - 4 n} \right) \\ s = \frac {1} {2} \left (m + \sqrt {m^2 - 4 n} \right) \\ r + s = m \\ r s = n \\ \] By the problem I posted before this one, we have \[ r^p + s^p \] is an integer say q, \[\left(m-\sqrt{m^2-4 n}\right)^p+\left(\sqrt{m^2-4 n}+m\right)^p=2^p \left(r^p+s^p\right)==2^p q \] We are done

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

Ahh yes it nicely packs the previous work xD I see now we may choose arbitrary m, n and a generate whole set of fun divisibility problems!

OpenStudy (phi):

If I did this correctly, we can show it's also divisible by 2^(2p) \[ \left( m-\sqrt{m^2-4 n} \right)^p+\left( \sqrt{m^2-4 n}+m \right)^p \\ \frac{\left(m+\sqrt{m^2-4 n}\right)^p}{\left(m+\sqrt{m^2-4 n}\right)^p}\left(m-\sqrt{m^2-4 n}\right)^p+\frac{\left(m-\sqrt{m^2-4 n}\right)^p}{\left(m-\sqrt{m^2-4n}\right)^p}\left( m+\sqrt{m^2-4 n}\right)^p \\ \frac{\left(m^2-m^2+ 4 n\right)^p}{\left(m+\sqrt{m^2-4 n}\right)^p}+\frac{\left(m^2-m^2+ 4 n\right)^p}{\left(m-\sqrt{m^2-4n}\right)^p} \\ \frac{2^{2p}n^p}{\left(m+\sqrt{m^2-4 n}\right)^p} + \frac{2^{2p}n^p}{\left(m-\sqrt{m^2-4 n}\right)^p} \\ 2^{2p}\left(\frac{n^p}{\left(m+\sqrt{m^2-4 n}\right)^p} + \frac{n^p}{\left(m-\sqrt{m^2-4 n}\right)^p} \right) \\ \]